OpenStudy (anonymous):

evaluate: integral 1/(x^2 sqrt(25+9x^2)) dx

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ x ^{2}\sqrt{25+9x ^{2}} }dx\]

OpenStudy (nikvist):